# MM-ACT: Learn from Multimodal Parallel Generation to Act

### v1

@@ 1 Introduction
Other works [18, 35, 13, 61] further extend the conventional state-to-action imitation learning paradigm into a visual-prediction-driven decision and planning framework. By incorporating visual prediction into the policy learning process, these approaches enable models to explicitly or implicitly model future visual dynamics, thereby achieving stronger predictability and planning capabilities in complex interactive environments. Although these world models [4, 37] excel in temporal and environmental dynamics, they are primarily trained for predictive objectives rather than task-oriented planning. This leads to limited instruction understanding and sub-task planning capabilities.
@@ 2.3 Vision-Language-Action Models
The Vision-Language-Action (VLA) model primarily aims to transform general visual-language inputs into a sequence of executable action outputs within a unified framework. Early methods built upon large-scale pre-trained Vision-Language Models (VLMs) by incorporating action heads or specialized action expert modules [5, 62, 20, 21, 3, 29, 22]. This allowed the model to simultaneously possess visual perception, language understanding, and action decision-making capabilities—achieving an end-to-end integration of vision, language, and action in one unified system. To bridge the latency gap between VLMs and the real-time demands of action systems, a dual-system design [24, 2, 32, 50, 30] is commonly adopted, decoupling the overall architecture into an upstream VLM and a downstream action-prediction head. The latter either directly decodes actions from the latent action tokens generated by the VLM, or employs an additional diffusion-based action head to the VLM’s outputs or intermediate hidden representations. Various reasoning strategies [49, 47, 8, 11, 60] have also been used to augment the action model, ranging from purely language-based reasoning to multimodal reasoning capabilities. However, these approaches typically perform diffusion-based fine-tuning for action generation directly on top of autoregressive (AR) pre-trained backbones, resulting in an objective misalignment between AR-based pre-training (token prediction) and diffusion-based fine-tuning (denoising). This inconsistency can introduce optimization misalignment and hinder the model’s ability to effectively leverage its pre-trained knowledge.
@@ 3.2 Context‑Shared Multimodal Learning
Figure 3: Training pipeline of $M$-ACT. Within the shared context, ground truth of three modalities are masked according to the decoding strategy of each modality, and appended to the context. The model take these inputs to perform forward processes across three modality generation tasks, computing the loss specifically on the masked tokens.
@@
Model Spatial SR (%) Object SR (%) Goal SR (%) Long SR (%) Average SR (%) OpenVLA [20] 84.7 88.4 79.2 53.7 76.5 $\pi_{0}$ + FAST [29] 96.4 96.9 98.8 90.6 91.5 $\pi_{0}$ [3] 96.9 96.2 98.2 92.0 92.0 OpenVLA-OFT [21] 96.2 98.3 96.2 90.7 95.4 CoT-VLA [57] 87.5 91.6 87.6 69.0 81.1 TraceVLA[58] 84.6 85.2 75.1 54.1 74.8 DreamVLA [54] 97.5 94.0 89.5 89.5 92.6 WorldVLA (512*512) [6] 87.6 96.2 83.4 60.0 81.8 UniVLA [40] 95.4 98.9 93.6 94.0 95.5 MM-ACT (Vanilla) 97.8 99.4 94.8 88.0 95.0 MM-ACT (+Text in Long) - - - 93.0(+5.0%) 96.3
@@ 4 Experiments
Model Adjust Bottle Beat Hammer Block Click Bell Dump Bin Bigbin $\pi_{0}$ [3] 89% 68% 40% 61% OpenVLA-OFT [21] 64% 7% 24% 31% MM-ACT (Vanilla) 51% 61% 86% 13% MM-ACT (+Text) 75% 67% 91% 13% MM-ACT (+Image) 72% 64% 91% 8% MM-ACT (+Text&Image) 71% 78% 95% 13% Model Moving Playingcard Away Lift Pot Place Burger Fries Place Can Basket $\pi_{0}$ [3] 47% 29% 41% 10% OpenVLA-OFT [21] 17% 3% 31% 8% MM-ACT (Vanilla) 30% 40% 46% 18% MM-ACT (+Text) 24% 28% 56% 18% MM-ACT (+Image) 39% 31% 72% 13% MM-ACT (+Text&Image) 39% 31% 73% 19% Overall Avg:     $\pi_{0}$ = 48.13%,      OpenVLA-OFT = 23.13%,      MM-ACT (Vanilla) = 43.13% MM-ACT (+Text) = 46.5%(+3.37%)     MM-ACT (+Image) = 48.75%(+5.62%)        MM-ACT (+Text&Image) = 52.38%(+9.25%)
@@ 4.3 Evaluation Results
As shown in Table 1, MM-ACT achieves the best average success rate of 96.3%, surpassing all existing baselines. Specifically, the vanilla version of MM-ACT achieves success rates of 97.8%, 99.4%, 94.8%, and 88.0% on Libero-Spatial, Libero-Object, Libero-Goal, and Libero-Long respectively. For VLM-based VLA, our approach outperforms OpenVLA by 19.8%, $\pi_{0}$ + FAST by 4.8%, $\pi_{0}$ by 4.3%, and OpenVLA-OFT by 0.9%. For Visual Prediction VLA, our method surpasses CoT-VLA by 15.2%, TraceVLA by 21.5%, and DreamVLA by 3.7%. In the case of Unified VLA, our approach exceeds WorldVLA by 14.5% and UniVLA by 0.8%. This validates the effectiveness of MM-ACT in terms of model architecture and training paradigms. Furthermore, as shown in the last two rows of the table, jointly optimizing the model for both task planning and action generation during training effectively improves its capability in long-horizon planning. In particular, the success rate on Libero-Long increases from 88.0% to 93.0%, achieving a notable improvement of +5.0%.
@@ 4.5 Ablation Study
$M$-ACT Overall Avg Time
one step PD,cs=8 43.13% 0.22s
re-mask PD,cs=8,t=6 42.38%(-0.75%) 1.06s
one step PD,cs=16 43.75% 0.23s
re-mask PD,cs=16,t=6 56.75%(+13.00%) 1.06s
@@ Appendix D Robotic Embodiments in Simulation & Real-World
Figure 10: Visualization of the Eight Tasks in RoboTwin2.0.
@@
Task: Moving Playingcard Away

### v2

@@ 1 Introduction
Other works [19, 37, 14, 63, 7] further extend the conventional state-to-action imitation learning paradigm into a visual-prediction-driven decision and planning framework. By incorporating visual prediction into the policy learning process, these approaches enable models to explicitly or implicitly model future visual dynamics, thereby achieving stronger predictability and planning capabilities in complex interactive environments. Although these world models [4, 39] excel in temporal and environmental dynamics, they are primarily trained for predictive objectives rather than task-oriented planning. This leads to limited instruction understanding and sub-task planning capabilities.
@@ 2.3 Vision-Language-Action Models
The Vision-Language-Action (VLA) model primarily aims to transform general visual-language inputs into a sequence of executable action outputs within a unified framework. Early methods built upon large-scale pre-trained Vision-Language Models (VLMs) by incorporating action heads or specialized action expert modules [5, 64, 21, 22, 3, 30, 23, 35]. This allowed the model to simultaneously possess visual perception, language understanding, and action decision-making capabilities—achieving an end-to-end integration of vision, language, and action in one unified system. To bridge the latency gap between VLMs and the real-time demands of action systems, a dual-system design [25, 2, 33, 52, 31] is commonly adopted, decoupling the overall architecture into an upstream VLM and a downstream action-prediction head. The latter either directly decodes actions from the latent action tokens generated by the VLM, or employs an additional diffusion-based action head to the VLM’s outputs or intermediate hidden representations. Various reasoning strategies [51, 49, 9, 12, 62] have also been used to augment the action model, ranging from purely language-based reasoning to multimodal reasoning capabilities. However, these approaches typically perform diffusion-based fine-tuning for action generation directly on top of autoregressive (AR) pre-trained backbones, resulting in an objective misalignment between AR-based pre-training (token prediction) and diffusion-based fine-tuning (denoising). This inconsistency can introduce optimization misalignment and hinder the model’s ability to effectively leverage its pre-trained knowledge.
@@ 3.2 Context‑Shared Multimodal Learning
Figure 3: Training pipeline of $M$-ACT. Within the shared context, ground truth of three modalities are masked according to the decoding strategy of each modality, and appended to the context. The model takes these inputs to perform forward processes across three modality generation tasks, computing the loss specifically on the masked tokens.
@@
Model Spatial SR (%) Object SR (%) Goal SR (%) Long SR (%) Average SR (%) OpenVLA [21] 84.7 88.4 79.2 53.7 76.5 $\pi_{0}$ + FAST [30] 96.4 96.8 88.6 60.2 85.5 $\pi_{0}$ [3] 96.8 98.8 95.8 85.2 94.2 OpenVLA-OFT [22] 96.2 98.3 96.2 90.7 95.4 CoT-VLA [59] 87.5 91.6 87.6 69.0 81.1 TraceVLA[60] 84.6 85.2 75.1 54.1 74.8 DreamVLA [56] 97.5 94.0 89.5 89.5 92.6 WorldVLA (512*512) [6] 87.6 96.2 83.4 60.0 81.8 UniVLA [42] 95.4 98.8 93.6 94.0 95.5 MM-ACT (Vanilla) 97.8 99.4 94.8 88.0 95.0 MM-ACT (+Text in Long) - - - 93.0(+5.0%) 96.3
@@ 4 Experiments
Model Adjust Bottle Beat Hammer Block Click Bell Dump Bin Bigbin $\pi_{0}$ [3] 89% 68% 40% 61% OpenVLA-OFT [22] 64% 7% 24% 31% MM-ACT (Vanilla) 51% 61% 86% 13% MM-ACT (+Text) 75% 67% 91% 13% MM-ACT (+Image) 72% 64% 91% 8% MM-ACT (+Text&Image) 71% 78% 95% 13% Model Move Playingcard Away Lift Pot Place Burger Fries Place Can Basket $\pi_{0}$ [3] 47% 29% 41% 10% OpenVLA-OFT [22] 17% 3% 31% 8% MM-ACT (Vanilla) 30% 40% 46% 18% MM-ACT (+Text) 24% 28% 56% 18% MM-ACT (+Image) 39% 31% 72% 13% MM-ACT (+Text&Image) 39% 31% 73% 19% Overall Avg:     $\pi_{0}$ = 48.13%,      OpenVLA-OFT = 23.13%,      MM-ACT (Vanilla) = 43.13% MM-ACT (+Text) = 46.5%(+3.37%)     MM-ACT (+Image) = 48.75%(+5.62%)        MM-ACT (+Text&Image) = 52.38%(+9.25%)
@@ 4.3 Evaluation Results
As shown in Table 1, MM-ACT achieves the best average success rate of 96.3%, surpassing all existing baselines. Specifically, the vanilla version of MM-ACT achieves success rates of 97.8%, 99.4%, 94.8%, and 88.0% on Libero-Spatial, Libero-Object, Libero-Goal, and Libero-Long respectively. For VLM-based VLA, our approach outperforms OpenVLA by 19.8%, $\pi_{0}$ + FAST by 10.8%, $\pi_{0}$ by 2.1%, and OpenVLA-OFT by 0.9%. For Visual Prediction VLA, our method surpasses CoT-VLA by 15.2%, TraceVLA by 21.5%, and DreamVLA by 3.7%. In the case of Unified VLA, our approach exceeds WorldVLA by 14.5% and UniVLA by 0.8%. This validates the effectiveness of MM-ACT in terms of model architecture and training paradigms. Furthermore, as shown in the last two rows of the table, jointly optimizing the model for both task planning and action generation during training effectively improves its capability in long-horizon planning. In particular, the success rate on Libero-Long increases from 88.0% to 93.0%, achieving a notable improvement of +5.0%.
@@ 4.5 Ablation Study
$M$-ACT Overall Avg (%) Time
one step PD, cs=8, t=1 43.13 0.22s
re-mask PD, cs=8, t=6 42.38(-0.75%) 1.06s
one step PD, cs=16, t=1 43.75 0.23s
re-mask PD, cs=16, t=6 56.75(+13.00%) 1.06s
@@ Appendix D Robotic Embodiments in Simulation & Real-World
Figure 10: Visualization of eight tasks in RoboTwin2.0.
@@
Task: Move Playingcard Away